2) Find the exact production elasticity of the production function Y= 10X - X^2, when X= 5.
Answer #1

The production function is given as:

Y = 10X - X​​​​​​2

The elasticity can be calculated by differentiating the production function with respect to X.

Differentiating with respect to X:

dY/dX = 10 - 2X

It is given that X = 5.

Substituting the value of X:

dY/dX = 10 - 2(5) = 10 - 10 = 0.

So, the production elasticity is equal to 0.